OpenStudy (anonymous):

find the total surface area of a cube with an edge of 11 ft. please show all work.

OpenStudy (anonymous):

726

OpenStudy (anonymous):

726 first you find one side 11 * 11 because its a cube and all the edges have the same length then you multiply that by the number of sides which is 6 so 11*11=121 121*6 = 726

OpenStudy (anonymous):

please explain.

OpenStudy (anonymous):

thanks.

OpenStudy (anonymous):

np

OpenStudy (anonymous):

The total surface area of a cube is equal to 6x^2, where x is the edge or lateral of a side and the symbol in between x and 2 means the power of. We can deduce this formula because a cube has 6 sides and the total surface area of a cube is the addition of the areas of each side. The area of one side is equal to x^2 or 11^2 = 11x11= 121. To find the surface area multiply the area by six to get a final answer of 726.

OpenStudy (anonymous):

726